spring of force constant k is stretched certain in distance.it takes twice as much work to strech second spring by half distance.force constant of second soring is

Grade:12th pass

1 Answers

erra akhil
8 years ago
Dear Anil,
work done {by a string streched by distance x} = (kx^2)/2. ----- eq.1
2 [work done {by a string stretched by distance x/2}] = (K(x/2)^2)/2).-------eq.2
On solving both equations we get,
K=8k.
k=1st spring constant.
K=2nd spring constant.
Approve if my answer helped you.